OpenStudy (anonymous):

Please help. Factor each completely. 1) 3x^2+5x-28 2) 81a^2-126a+49

OpenStudy (lgbasallote):

(3x-7)(x+4) and (9a-7)^2

OpenStudy (anonymous):

is that the whole process? i dont understand how you got to that point. Thank you

sam (.sam.):

3x^(2)+5x-28 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (-84) that add up to b (5).In this problem 4*-(7)/(3)=-(28)/(3) (which is (c)/(a)) and 4-(7)/(3)=(5)/(3) (which is (b)/(a) , so insert 4 as the right hand term of one factor and -(7)/(3) as the right-hand term of the other factor. (x+4)(x-(7)/(3)) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (x+4)(3x-7)

OpenStudy (anonymous):

great i think i got it sounds kinda hard